# Meta-DAN: towards an efficient prediction strategy for page-level handwritten text recognition

### v1

@@ Abstract
Recent advances in text recognition led to a paradigm shift for page-level recognition, from multi-step segmentation-based approaches to end-to-end attention-based ones. However, the naïve character-level autoregressive decoding process results in long prediction times: it requires several seconds to process a single page image on a modern GPU. We propose the Meta Document Attention Network (Meta-DAN) as a novel decoding strategy to reduce the prediction time while enabling a better context modeling. It relies on two main components: windowed queries, to process several transformer queries altogether, enlarging the context modeling with near future; and multi-token predictions, whose goal is to predict several tokens per query instead of only the next one. We evaluate the proposed approach on 10 full-page handwritten datasets and demonstrate state-of-the-art results on average in terms of character error rate. Source code and weights of trained models are available at https://github.com/FactoDeepLearning/meta_dan.
Index Terms:
@@
I Introduction
@@
Handwritten Text Recognition (HTR) is the process of automatically extracting handwritten text from a digitized image and converting it into machine-readable text. The historical approach to tackle this task, still in the majority today, consists of a multi-stage paradigm. A document image is segmented into sections of text, which are then ordered so as to preserve a coherent reading order. Thereafter, these text sections are recognized and reassembled to obtain the overall transcription of the document. Over time, the nature of the textual section has broadened, from the isolated character [1, 2], to the word [3, 4], then to the line [5, 6, 7] and more recently to the paragraph [8, 9, 10, 11, 12], gradually increasing the context usable during recognition. However, regardless of the granularity used, this multi-step approach comes with significant drawbacks: 1) the segmentation step relies on a neural network that needs to be trained, requiring additional segmentation annotations, 2) the ordering step is generally a hand-crafted algorithm that needs to be adapted to the layout of each target dataset, 3) the recognition of the individual sections is carried out independently, limiting the context modeling during prediction, 4) errors accumulate in each of these steps. Also, it has to be noted that works of the literature generally focus on the segmentation or recognition task separately, so that the given recognition results are based on a human-generated segmentation, which is not representative of the task in real-life conditions.
@@
More recently, the Document Attention Network (DAN) [13] was proposed as the first end-to-end approach for HTR at page level, achieving similar recognition performance compared to the multi-step approach. In addition to text, the DAN also predicts special XML-like layout tokens, leading to enriched serialized representations of documents. The DAN architecture includes a transformer-based decoder to perform a character-level autoregressive prediction process. Indeed, the transformer’s attention mechanism is used to generate an implicit segmentation of characters, one after the other, making the model learn the reading order of the document through a text-only supervision. Due to its state-of-the-art performance, the DAN is currently the default approach used for end-to-end handwritten text recognition at page-level [14, 15]. In addition, it has been extended to handle other tasks at page level such as named entity recognition [16, 17, 18], table recognition [19], and key-value information extraction [20, 21, 22], by adding dedicated tokens in the predictable token set.
@@
The main drawback of the DAN, which is also an advantage, lies in its character-level iterative prediction process, which is slow when processing whole pages. Indeed, on the one hand, processing text at character level reduces the vocabulary used, making training more easy. It also makes it possible to manipulate lower-dimensional latent states, since each token carries little information, reducing computational complexity. On the other hand, it increases the size of the generated sequences, and thus, the number of iterations to be performed and the prediction time.
@@
In this work, we propose the Meta Document Attention Network (Meta-DAN), a novel prediction strategy to improve both inference speed and recognition accuracy for end-to-end page-level handwritten text recognition. It lies on the combined use of two distinct approaches. The first one is based on multi-token predictions as it showed interesting results for Natural Language Processing (NLP) [23]. The second consists in windowed queries, i.e., processing successive transformer queries altogether as a sliding window to process them in parallel.
The main contributions of our work are summarized as follows:
@@
  • We study the introduction of a specific dynamic prediction policy to strike a balance between prediction speed and recognition performance for multi-token predictions.
@@
  • We provide an extensive evaluation over a set of 10 datasets, covering several languages and alphabets. It highlights that combining both proposed approaches (Meta-DAN) leads to on-average improved recognition performance compared to concurrent approaches.
  • We show that multilingual page-level text recognition can be beneficial to increase the recognition performance for the mainly represented languages.
II Related Works
II-A Handwritten text recognition
Over the last decades, the field of handwriting recognition has made major advances, moving from isolated character recognition based on expert models, to end-to-end recognition of full-page documents. These advances are mainly due to the modeling capability of deep neural networks, which represent the state of the art in this field. The approaches proposed in the literature can be divided into two main categories we refer to as the alignment-based and the attention-based approaches.
The alignment-based approaches consider the recognition task as a one-dimensional sequence alignment problem between a prediction sequence and a ground truth sequence. Modeling the text recognition task in this way enables the end-to-end recognition of word [24, 25] and line [5, 7] images. The practical application of this method is based on two aspects: the reduction of the vertical dimension, generally achieved through convolution or pooling, and the use of the Connectionist Temporal Classification (CTC) [26] introduced in 2006. The CTC loss enables to align a probability lattice computed from the input image with a sequence of characters corresponding to the ground truth. The strength of this approach is also its weakness: it only deals with one-dimensional sequences. To go a step further, two works showed that it is feasible to recognize whole paragraphs by using tricks to assimilate a paragraph to a long line of text, through bilinear interpolations [10], or concatenation [27]. However, to the best of our knowledge, this alignment-based approach is not suited for end-to-end full-page recognition, since the complexity of page layouts (e.g., multi-column pages) makes it impossible to reduce it to a one-dimensional sequence alignment problem.
@@ II-A Handwritten text recognition
More recently, the attention-based approaches emerged, particularly with the rise in popularity of the Transformer architecture [28]. They rely on attention mechanisms to provide an iterative decoding process at character level, performing a kind of implicit character segmentation through the attention maps. In other words, characters are predicted one after the other following an autoregressive process. At each decoding step, attention weights are computed and matched to the extracted features of the input image, with the aim of focusing on a specific part of the image (the character to recognize). Contrary to the alignment-based approaches, the attention-based approaches can be applied to any input image, from a single line [6, 29] or paragraph [9, 30, 11] to a two-page document with the DAN-based works [13, 14, 17]. The main drawback of such approach lies in the prediction time that grows with the length of the text sequence to be recognized.
@@
It has to be noted that marginal works [8, 12] were proposed to handle paragraph recognition with a hybrid approach combining the attention mechanisms, to generate an implicit line segmentation, and the alignment-based approach, to recognize them iteratively through CTC decoding. In addition, an attempt of adapting the CTC to 2D-to-1D alignment was proposed in [31], but it did not lead to convincing results.
@@
In summary, alignment-based approaches are advantageous because of their one-shot prediction, but they are limited to images with a single column of text, whereas attention-based approaches are capable of handling entire pages with complex layout but involve longer prediction times. Attention-based decoding strategies also suffer from a limited context modeling, in that it only has access to the previously predicted tokens to make its prediction (only past is considered). In this work, we focus on reducing the prediction time of the attention-based character-level approach to reach fast end-to-end page-level predictions, while improving the context modeling through the integration of near future in the decoding process.
II-B Speeding up autoregressive processes
The prediction time issue of autoregressive processes can be found in various applications including a text generation stage, and several strategies have been proposed. In [32], a transformer-based network is trained to predict the $n$ future tokens instead of a single one in the context of text summarization. We refer to this strategy as multi-token prediction. This idea was successfully applied for large language models recently [23]. In this latter work, $k$ prediction heads are applied on the same transformer output, each one predicting one of the $k$ next tokens, limiting the computational overhead. A similar approach was proposed for neural machine translation in [33] by including an extra verification step: prediction is performed another time using all the predictions as in teacher forcing to validate them or not.
The authors of [34] proposed another way to reduce the prediction time for the task of image captioning. The target sequence is split into subgroups of words, which are decoded in parallel. They use an extra DEL token to remove duplicated predictions. Closer to our work, a similar approach was proposed to speed up the Document Attention Network by splitting the decoding process into two stages [14]. The first stage is dedicated to predicting the first character of each line, and the second one consists in completing all lines in parallel, reducing the number of decoding steps. While the prediction time is significantly decreased, this alternative suffers from a quite chaotic context: the model has access to the beginning of each line, i.e., a partial view of both past and future. The windowed queries we propose also rely on the idea of processing queries in parallel, but they differ in that the processed queries are successive, leading to a unique decoding stage and a context mixing full past and near future.
An alternative to speed up the prediction process is to reduce the length of the target sequence. This can be achieved through a specific tokenizer, one of the best known is the Byte-Pair Encoding (BPE) algorithm [35]. Instead of encoding the text to be recognized as a sequence of character tokens, it is encoded as tokens representing sub-words of variable sizes, obtained from their frequency of occurrence in a given dataset. This technique has been applied in [17] showing a significant prediction time decrease, in return for a higher character error rate. The approach we propose is compatible with this kind of technique as it rely on the token prediction strategy, regardless of the nature of the tokens.
@@ II-B Speeding up autoregressive processes
In this paper, we propose to combine both windowed queries and multi-token predictions to speed up the prediction process while improving the context modeling capability of the model. While evaluated for the HTR task in this paper, the approach is generic and could be applied for any task relying on a transformer-based iterative decoding process.
III Background
DAN architecture. Encoding stage: 2D features $\boldsymbol{f}$ are extracted from the input image through a Fully Convolutional Network (FCN). Features are enhanced with 2D positional encoding and flattened for transformer needs. Decoding stage: at each iteration $t$, a query sequence $\boldsymbol{q}^{t}$ is generated from the previous predictions $\hat{\boldsymbol{y}}^{t}$ and associated to the image features through transformer’s attention mechanisms to compute the new prediction $\hat{\boldsymbol{y}}_{t}$.
To make things clearer, we first propose a formulation of the handwritten text recognition problem, we go over the main points of the DAN [13] approach on which is built this work, and we discuss the Faster DAN [14] proposition.
III-A Problem formulation
@@
III-B Document Attention Network
@@
The Document Attention Network (DAN) [13] was proposed to tackle this task as a character-level autoregressive prediction process. While the original approach generates additional layout tokens in addition to text tokens, we only focus on text recognition in this work. The DAN refers to a deep neural network illustrated in Figure III. It is made up of two main building blocks: a Fully Convolutional Network (FCN) encoder and a causal transformed-based decoder. The general idea of the approach is as follows. Based on the image features extracted by the encoder, and on the previously predicted tokens, the goal is to predict the next token (each token representing a character).
@@
The DAN encoder, noted $e$, was first proposed in [12] as a generic feature extractor for handwritten text recognition able to deal with images of variable sizes. It mainly consists in a succession of 18 standard convolutions and 12 Depthwise Separable Convolutions [36], coupled with instance normalization, mix dropout layers and skip connections. It extracts 2D features $\boldsymbol{f}=e(\boldsymbol{X})$ ($\boldsymbol{f}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$) from the input image by reducing the height by 32 and the width by 8 ($H_{f}=\frac{H}{32}$, $W_{f}=\frac{W}{8}$ and $C_{f}=256)$. All details can be found in the original paper [12].

### v2

@@ Abstract
Recent advances in text recognition led to a paradigm shift for page-level recognition, from multi-step segmentation-based approaches to end-to-end attention-based ones. However, the naïve character-level autoregressive decoding process results in long prediction times: it requires several seconds to process a single page image on a modern GPU. We propose the Meta Document Attention Network (Meta-DAN) as a novel decoding strategy to reduce the prediction time while enabling better context modeling. It relies on two main components: windowed queries, to process several transformer queries altogether, enlarging the context modeling with the near future; and multi-token predictions, whose goal is to predict several tokens per query instead of only the next one. We evaluate the proposed approach on 10 full-page handwritten datasets and demonstrate state-of-the-art results on average in terms of character error rate. Source code and weights of trained models are available at https://github.com/FactoDeepLearning/meta_dan.
keywords:
@@
journal: Pattern Recognition
[…]
organization=Univ Rennes, CNRS, IRISA - UMR 6074, F-35000 Rennes,country=France
1 Introduction
@@
Handwritten Text Recognition (HTR) is the process of automatically extracting handwritten text from a digitized image and converting it into machine-readable text. The historical approach to tackle this task, still in the majority today, consists of a multi-stage paradigm. A document image is segmented into sections of text, which are then ordered to preserve a coherent reading order. Thereafter, these text sections are recognized and reassembled to obtain the overall transcription. Over time, the nature of the textual section has broadened, from the isolated character [1], to the word [21], then to the line [45, 32, 12] and more recently to the paragraph [4, 3, 48, 15], gradually increasing the context usable during recognition. However, regardless of the granularity, this multi-step approach comes with significant drawbacks: 1) the segmentation step requires an additional segmentation model and annotations, 2) the ordering step is generally a dataset-specific hand-crafted algorithm, 3) the recognition of the individual sections is carried out independently, limiting the context modeling during prediction, 4) errors accumulate in each of these steps.
@@
More recently, the Document Attention Network (DAN) [14] was proposed as the first end-to-end approach for HTR at page level, achieving similar recognition performance compared to the multi-step approach. The DAN architecture includes a transformer-based decoder to perform a character-level autoregressive prediction process. Due to its state-of-the-art performance, the DAN is currently the default approach used for end-to-end handwritten text recognition at page-level [16, 8]. In addition, it has been extended to handle other tasks at page level such as named entity recognition [11, 10], table recognition [5], and key-value information extraction [41, 40], by adding dedicated tokens in the predictable token set.
@@
The main drawback of the DAN, which is also an advantage, lies in its character-level iterative prediction process. Indeed, on the one hand, processing text at character level reduces the vocabulary, making training easier. It also enables to manipulate lower-dimensional latent states, since each token carries little information, reducing computational complexity. On the other hand, it increases the target sequence length, and so the number of decoding iterations, increasing the prediction time.
@@
In this work, we propose the Meta Document Attention Network (Meta-DAN), a novel prediction strategy to improve both inference speed and recognition accuracy for end-to-end page-level handwritten text recognition. It lies on the combined use of two distinct approaches. The first one is based on multi-token predictions as it showed interesting results for Natural Language Processing (NLP) [23]. The second consists of windowed queries, i.e., processing successive transformer queries altogether as a sliding window to process them in parallel. Our main contributions are as follows:
@@
  • We study the introduction of a confidence-based dynamic prediction policy to strike a balance between prediction speed and recognition performance for multi-token predictions.
@@
  • We provide an extensive evaluation over 10 various public datasets. To our knowledge, this is the largest benchmark evaluated at page-level. This evaluation highlights that combining both proposed approaches (Meta-DAN) leads to on-average state-of-the-art recognition performance.
2 Related Works
2.1 Handwritten text recognition
State-of-the-art approaches can be divided into two main categories we refer to as the alignment-based and the attention-based approaches.
The alignment-based approaches consider the recognition task as a one-dimensional sequence alignment problem between a predicted probability lattice and a ground truth sequence. The Connectionist Temporal Classification (CTC) [24] was proposed in 2006 to perform this alignment, enabling the end-to-end recognition of word [25, 7] and line [45, 12] images.
However, the CTC only deals with one-dimensional sequences. This way, tricks have been proposed to assimilate a paragraph to a long line of text, through bilinear interpolations [48], or concatenation [13]. But, to our knowledge, alignment-based approaches are not suited for end-to-end full-page recognition, since the complexity of page layouts (e.g., multi-column pages) prevent to reduce it to a one-dimensional sequence alignment problem.
@@ 2.1 Handwritten text recognition
More recently, the attention-based approaches emerged, particularly with the rise in popularity of the Transformer architecture [43]. They rely on attention mechanisms to provide an iterative autoregressive decoding process at character level, performing a kind of implicit character segmentation through the attention maps.
Contrary to the alignment-based approaches, the attention-based approaches can be applied to any input image, from a single line [32, 2] or paragraph [3, 38] to a two-page document with the DAN-based works [14, 16, 11].
@@
It has to be noted that marginal works [4, 15] were proposed to handle paragraph recognition with a hybrid approach combining the attention mechanisms, to generate an implicit line segmentation, and the alignment-based approach, to recognize them iteratively through CTC decoding. In addition, an attempt to adapt the CTC to 2D-to-1D alignment was proposed in [36], but it did not lead to convincing results.
@@
In summary, alignment-based approaches are advantageous because of their one-shot prediction, but they are limited to images with a single column of text, whereas attention-based approaches are capable of handling entire pages with complex layouts, but involve longer prediction times due to their autoregressive nature.
In this work, we focus on reducing the prediction time of the attention-based character-level approach to reach fast end-to-end page-level predictions, while improving the context modeling through the integration of the near future in the decoding process.
2.2 Speeding up autoregressive processes
The prediction time issue of autoregressive decoders is not limited to HTR. In [33], a multi-token prediction strategy was proposed to speed up text summarization: a transformer model is trained to predict the $n$ future tokens at once using $n$ projection heads. This idea was also successfully applied for neural machine translation [39] and for large language models [23].
The authors of [22] proposed another way to reduce the prediction time for the task of image captioning. The target sequence is split into subgroups of words, which are decoded in parallel.
Closer to our work, a similar approach was proposed to speed up the Document Attention Network by splitting the decoding process into two stages [16]. The first stage is dedicated to predicting the first character of each line, and the second one consists in completing all lines in parallel, reducing the number of decoding steps.
The windowed queries we propose also rely on the idea of processing queries in parallel, but they differ in that the processed queries are successive, leading to a unique decoding stage and a context mixing full past and near future.
An alternative to speed up the prediction process is to reduce the length of the target sequence, by relying on a dedicated tokenizer, such as the Byte-Pair Encoding (BPE) algorithm [37].
This technique was applied in [11], showing a significant prediction time decrease, in return for a higher character error rate. The approach we propose is compatible with this kind of technique as it relies on the token prediction strategy, regardless of the nature of the tokens.
@@ 2.2 Speeding up autoregressive processes
In this paper, we propose to combine both windowed queries and multi-token predictions to speed up the prediction process while improving the context modeling capability of the model.
3 Background
3.1 HTR problem formulation
@@
3.2 Document Attention Network
@@
The Document Attention Network (DAN) [14] was proposed to tackle this task as a character-level autoregressive prediction process. It refers to a deep neural network illustrated in Figure 3.2. It is made up of two main building blocks: a Fully Convolutional Network (FCN) encoder and a causal transformer-based decoder. The general idea of this approach is as follows. Based on the image features extracted by the encoder, and on the previously predicted tokens, the goal is to predict the next token (i.e., a character). To determine when to stop this iterative decoding process, a special end-of-transcription token (<e>) is added to the set of predictable tokens $\mathcal{A}$, and appended to the ground truth sequence: $\boldsymbol{y}\in\mathcal{A}^{L^{y}+1}$ with $\mathcal{A}=\mathcal{A}^{c}\cup\{\text{<e>}\}$ and $\boldsymbol{y}_{L^{y}+1}=\text{<e>}$.
DAN architecture. Encoding stage: 2D features $\boldsymbol{f}$ are extracted from the input image through a Fully Convolutional Network (FCN). Decoding stage: at each iteration $t$, a query sequence $\boldsymbol{q}^{t}$ is generated from the previous predictions $\hat{\boldsymbol{y}}^{t}$ and associated with the image features through the transformer’s attention mechanisms to predict the next token $\hat{\boldsymbol{y}}_{t}$. The decoding is initialized with a start-of-prediction token $\hat{\boldsymbol{y}}_{0}=\text{<s>}$ and stops when the end-of-transcription token <e> is predicted.
Regarding the notations, given that matrices change over the decoding iterations, we opted for the following convention. The exponent specifies the version of the matrix at a given time, and the index specifies a vector from that matrix. For instance, in Figure 3.2, $\boldsymbol{q}^{t}$ is the matrix of query token vectors at decoding iteration $t$ made up of $t$ vectors indexed from 0 to t1t-1. Given that the vector $\boldsymbol{q}_{i}^{t}$ at index $i$ is the same for all $t$, it can be referred to as $\boldsymbol{q}_{i}$, omitting the exponent for readability. In the following, we describe the different stages implied by the DAN by referring to the circled numbers indicated in Figure 3.2.
1
@@
The DAN encoder, noted $\text{Enc}_{\text{DAN}}$, extracts 2D features $\boldsymbol{f}=\text{Enc}_{\text{DAN}}(\boldsymbol{X})$ ($\boldsymbol{f}\in\mathbb{R}^{H_{f}\times W_{f}\times C_{f}}$) from the input image by reducing the height by 32 and the width by 8 $(H_{f}=\frac{H}{32}$, $W_{f}=\frac{W}{8}$ and $C_{f}=256)$. This encoder was first proposed in [15] as a generic feature extractor for handwritten text recognition able to deal with images of variable sizes. All details can be found in the original paper [15].
2 For transformer needs, positional information is injected into the image features $\boldsymbol{f}$ through additive 2D positional encoding $\boldsymbol{\pi}_{\text{2D}}$ [38]. Height and width are each encoded over half of the vector channels. The result is reshaped to a one-dimensional representation: $\boldsymbol{g}=\text{flatten}(\boldsymbol{f}+\boldsymbol{\pi}_{\text{2D}})$.
3 For each iteration $t$, the previously predicted tokens are put together to be used by the model to know which character token should be predicted next: $\hat{\boldsymbol{y}}^{t}=[\hat{\boldsymbol{y}}_{0},\dots,\hat{\boldsymbol{y}}_{t-1}]$. A special start-of-transcription token (<s>) is used to initiate the prediction process: $\hat{\boldsymbol{y}}_{0}=\text{<s>}$.
4 A trainable weight matrix $\boldsymbol{E}\in\mathbb{R}^{(|\mathcal{A}|+1)\times C_{f}}$ is used to map each token of the vocabulary $\mathcal{A}$ to a vector of dimension $C_{f}$, leading to $\boldsymbol{e}^{t}\in\mathbb{R}^{t\times C_{f}}$, with $\boldsymbol{e}^{t}_{i}=\boldsymbol{E}_{\hat{\boldsymbol{y}}_{i}}$. One can note the additional embedding vector (+1), which corresponds to the start-of-transcription token (<s>).
5 As the transformer attention mechanism is permutation equivariant, one-dimensional positional encoding $\boldsymbol{\pi}^{t}_{\text{1D}}$ [43] is added to these embeddings to preserve the order information of the tokens, resulting in the query sequence $\boldsymbol{q}^{t}=\boldsymbol{e}^{t}+\boldsymbol{\pi}^{t}_{\text{1D}}$.
6 The query tokens $\boldsymbol{q}^{t}$ are processed through the DAN decoder, noted $\text{Dec}_{\text{DAN}}$, which corresponds to 8 stacked transformer decoder layers preserving the number of channels $C_{f}$. The goal is to extract relevant information from the flattened image features $\boldsymbol{g}$, i.e., a latent representation $\boldsymbol{o}^{t}_{t-1}$ of the next character to predict: $\boldsymbol{o}^{t}=\text{Dec}_{\text{DAN}}(\boldsymbol{q}^{t},\boldsymbol{g})$.
7 Based on the last output of the decoder $\boldsymbol{o}_{t-1}$, a score vector $\boldsymbol{s}_{t}\in\mathbb{R}^{|\mathcal{A}|}$ is computed through a linear projection layer of trainable weights $\boldsymbol{W}\in\mathbb{R}^{|\mathcal{A}|\times C_{f}}$ to associate a score with each predictable token:
$\boldsymbol{s}_{t}=\boldsymbol{W}\boldsymbol{o}_{t-1}.$ (1)
8
The predicted token $\hat{\boldsymbol{y}}_{t}$ corresponds to the one with the highest score: $\hat{\boldsymbol{y}}_{t}=\operatorname*{arg\,max}(\boldsymbol{s}_{t})$.
Steps 3 to 8 are repeated until the specific end-of-transcription token (<e>) is predicted.
$\mathcal{L}_{\text{DAN}}=\sum_{t=1}^{L^{y}+1}\mathcal{L}_{\text{CE}}(\boldsymbol{y}_{t},\boldsymbol{p}_{t}),$ (2)
where $\boldsymbol{p}_{t,i}=\frac{e^{\boldsymbol{s}_{t,i}}}{\sum_{j}e^{\boldsymbol{s}_{t,j}}}$ is the predicted probability associated with token $i$ at time $t$.
3.3 Faster DAN
The Faster DAN [16] was proposed as a decoding alternative to speed up the inference stage. It aims to parallelize the text line predictions by relying on a two-stage approach. First, an autoregressive process predicts the first character of each line, following the original DAN strategy. Then, all lines are completed in parallel, reducing the total number of iterations. The main difference appears in that second decoding stage. Instead of using the output corresponding to the last token $\boldsymbol{o}_{t-1}$ only, the token prediction is performed using the outputs corresponding to the last token of each text line. In this way, for $N$ text lines, there are $N$ token predictions per iteration. Equation 1 becomes: $\boldsymbol{s}^{l}_{t}=\boldsymbol{W}\boldsymbol{o}^{l}_{t-1}$, where $\boldsymbol{s}^{l}_{t}$ is the score vector associated with the token prediction of the next character for line $l$ at iteration $t$, and $\boldsymbol{o}^{l}_{t-1}$ is the transformer output corresponding to the last predicted token for line $l$.
Although the Faster DAN showed interesting inference speed improvements, it results in a loss of past context when predicting a character. Indeed, instead of benefiting from the whole past context as is the case for the DAN, the Faster DAN has a partial view of the past and future due to the parallel line predictions, which reduces its ability to model the language.
The strategy we propose aims to get the best of both worlds: improving context modeling by processing successive queries altogether through multi-token predictions, while reducing prediction time through parallelization.
4 Meta Document Attention Network
We propose two new approaches for the autoregressive prediction of text tokens in the context of end-to-end handwritten text recognition at page level. For both, the idea is to predict multiple tokens at each decoding iteration to reduce the number of decoding steps, while preserving the whole past context. We assume this is a key element to reduce the number of recognition errors through language modeling. First, we propose to perform multiple predictions from the same query $\boldsymbol{q}_{t-1}$, to predict the $m$ next tokens at once. This is referred to as the Document Attention Network with Multi-Tokens predictions (MT-DAN). Second, we propose a query approach based on a window mechanism to process multiple queries at once, each associated with a single prediction. This is referred to as the Document Attention Network with Windowed queries (W-DAN). Finally, we design a global strategy including these two approaches, called Meta Document Attention Network (Meta-DAN). A visual comparison of the prediction processes is depicted in Figure 4.